### v3

@@ 4.1 Data
We collect 120 monthly variables for the US spanning the period January 1959 to December 2023. The data, which are obtained from the Federal Reserve Economic Data (FRED) and are available at https://fred.stlouisfed.org, cover a wide range of key macroeconomic variables that applied economists monitor regularly, such as different measures of output, prices, interest and exchange rates, and stock market performance. We provide a full list of the data and their transformations in order to achieve stationarity in Appendix B. Out of the 120 series, we further distinguish a subset of “variables of interest”, that is, key variables of interest which we will inspect very closely in order to evaluate how well the different models perform. These variables, along with their transformation codes, are described in Table 2.
@@ 4.2 Forecasting Exercise
We use the first twenty-five year of data, January 1960–December 1984, to obtain initial parameter estimates for all the econometric models, which are then used to predict outcomes from January 1985 ($h=1$) to December 1985 ($h=12$). The next period, we include data for January 1985 in the estimation sample, and use the resulting estimates to predict the outcomes from February 1985 to January 1986. We proceed recursively in this fashion until December 2022, thus generating a time series of forecasts for each forecast horizon $h$, with $h=1,...,12$. Note that when $h>1$, point forecasts are iterated and predictive simulation is used to produce the predictive densities.
Starting with the VARs, we estimate models of three different sizes: Medium (the 19 variables listed in Table 2), Large (the variables in medium plus an additional 20), and X-large (all 120 series available), that is, we consider 19, 39 and 120-variable sets. All models have a lag length of $p=6$. For the factor models, we follow the same strategy. Namely, for each variable size and each sample, we first extract the first common factor, $\hat{f}_{1t}$. Next, for each variable in that same set and each forecast horizon, we run the factor-augment autoregression in (13) and forecast up to $12$ months ahead. Similarly, for BART we re-estimate the trees for each variable set and sample, as the number of series in $X_{-i,t-h}$ changes with each variable set.
We proceed in a similar fashion with all TSLMs. More specifically, for each TSLM considered, starting with the first forecast date (December 1984), we feed the historical data to the pretrained model (when working with the univariate TSLMs we supplied one time series at a time, while with the multivariate TSLMs we supplied all time series within the specific model size at once). Next, we query the model to generate a zero-shot forecast for each of the variables provided, up to 12 steps ahead. [^14] For all our experiments, we downloaded the open-sourced pretrained models and run them locally on our computer, using an Nvidia A6000 GPU with 48GB of memory. The only exception was Time-GPT, where we instead queried directly their API, with a separate query for each forecast date-time series combination. We also include a benchmark approach which uses OLS forecasts from univariate AR(1) models.
@@
where $\underline{t}$ and $\overline{t}$ denote the start and end of the out-of-sample period, and where $e_{i,j,\tau+h}^{2}$ and $e_{bcmk,j,\tau+h}^{2}$ are the squared forecast errors of variable $j$ at time $\tau$ and forecast horizon $h$ associated with model $i$ ($i\in\left\{\text{BVAR v1},\text{BVAR v2},\text{Factor model},\text{BART}\right\}\cup{TSLM}{}s$) and the AR(1) model, respectively. The point forecasts used to compute the forecast errors are obtained by averaging over the draws from the various models’ $h$-step-ahead predictive densities. Values of $RMSFE_{ijh}$ below one suggest that model $i$ produces more accurate point forecasts than the AR(1) benchmark for variable $j$ and forecast horizon $h$.
In closing this Section it is worth mentioning that in this paper we are only focusing on point forecast accuracy. This choice is intentional. While there is a rich literature on density forecasts for macroeconomic variables, using TSLM to construct density forecasts presents additional challenges, and we leave that to further research.
@@ 5 Results
5.1 Do some LLMs Perform Better than Others?
To begin, we focus on the overall forecasting performance of the various TSLMs under examination. Given that the evaluation sample we are considering include the Covid-19 period, to avoid conflating the overall picture with the large idiosyncrasies brought about by the pandemic, we focus first on the accuracy of the various models by stopping before the onset of the pandemic. That is, our evaluation sample stops in December 2019. Later on, we analyze how stable these are results are over time, and also look at the accuracy of the various model over the most recent period, January 2020 to December 2023. Figure 2 provides a synoptic view of the forecasting performance, as measured by RMSFE ratios relative to the AR benchmark. Each box-plot represents the interquartile range of the distribution of relative RMSFEs. When the mass is concentrated above the line corresponding to $RMSFE=1$, this is an indication that the models do generally worse than the benchmark. The solid line within the box represents the median, while the whiskers represent the maximum and minimum data points that are not defined as outliers. [^15] Outliers are defined as values that are more than 1.5 times the interquantile range away from the top or bottom of the box (i.e. from the lower and upper quartiles).
@@ 5.1 Do some LLMs Perform Better than Others?
Outliers are not depicted for scaling reasons, and instead we decided to separately report the maximum and minimum RMSFEs attained by each model in Table 3 (together with their median and standard deviation). Indeed, it does happen on several occasions that for some series the TSLMs produce forecasts that are patently unreasonable. These forecasts would be automatically disregarded by the user and would call for a fine-tuning of the models. This result highlights that while zero-shot TSLMs may produce on average reasonable forecasts, they should not be used mechanically and do require careful monitoring.
Having said that, the pattern emerging from a careful inspection of Figure 2 is one of heterogeneity among the various TSLMs. Specifically, some TSLMs show a performance that is consistently worse than others. These are TTM-enhanced and Time-GPT, for which the box plots are invariably positioned higher than those of the remaining TSLMs, signaling higher RMSFEs. These models also perform poorly compared to the benchmark, for example Time-GPT has large part of the mass of the box-plot above 1 for all horizons, and TTM for the one-step-ahead horizon. Lag-LLama performs a bit better than these two, however it is also systematically outperformed by the remaining models.
The TSLMs showing the best forecasting performance are Moirai and TimesFM. As we discussed, Moirai comes in three sizes, and the evidence shows that the largest size performs the best, with a mass of RMSFEs being consistently below 1 at all horizons. Because of these considerations, in the remainder of this Section, we will only focus on these best performing TSLMs (Moirai Large and TimesFM), and compare their performance against the BVARs and the factor model.
5.2 Differences in Point Forecast between LLMs and Econometric Models
@@
Focusing first on the econometric models, it is remarkable that both the BVAR with natural conjugate prior and the factor model are consistently showing a good performance relative to the AR model. This is a well known result, which has been obtained by several different studies using this dataset. Surprisingly to us, the BVAR with asymmetric priors performs slightly worse than the symmetric counterpart, a result that is at odds with previous findings.
As already shown in Figure 2, Moirai Large and TimesFM also perform consistently better than the AR benchmark, but the distribution of RMSFEs is similar to that obtained with the econometric models. If anything, while all the models at almost all horizons have a median that is below the threshold of 1, there is a tendency for the distribution of the RMSFEs from TSLMs to have a more pronounced right tail, signaling an increased likelihood of observing large forecast errors. Instead, the econometric methods produce RMSFE distributions that are more skewed to the left, signaling a consistent better performance compared to the AR.
As before, additional descriptive statistics including the maximum and minimum relative RMSFEs (including outliers) are reported in Table 4. From this table it is clear that the econometric models do not present any extreme cases in which forecast errors are a multiple of those obtained by the AR benchmark, which it is instead the case for the TSLMs results shown in Table 3. Also, it is evident that the RMFSEs coming out of the econometric models are less dispersed. Overall the picture that emerges is one in which BVARs and factor models seem to offer a more stable and reliable performance, when compared to zero-shot TSLMs.
@@
5.3 Are the Differences Statistically Significant?
The previous subsections have considered the distribution of RMSFEs relative to an AR benchmark where the distribution was computed across all of the series in the dataset. The fact that such distribution was to the left (right) of the threshold of 1 meant that overall a model was producing better (worse) forecasts than the AR model, but no assessment was made on whether each forecast individually was statistically different from the one produced by the benchmark. In this subsection we tackle this specific issue. To save space, we only consider the 19 variables included in the medium model. A full set of result, covering all 120 variables, is available upon request.
To provide a rough gauge of whether the RMSE ratios are significantly different from 1, we use the Diebold and Mariano (1995) t-statistic (DM). Our use of the DM test with models that are in some case nested is a deliberate choice. Monte Carlo evidence in Clark and McCracken (2012, 2015) indicates that, with nested models, the Diebold-Mariano test compared against normal critical values can be viewed as a somewhat conservative (conservative in the sense of tending to have size modestly below nominal size) test for equal accuracy in the finite sample. Differences that are statistically significant are denoted by one, two, or three stars, corresponding to significance levels of 10%, 5%, and 1%, respectively. The underlying p-values are based on t-statistics computed with a serial correlation-robust variance, using a rectangular kernel, $h-1$ lags, and the small-sample adjustment of Harvey et al. (1997).
Results for forecast horizons of one and three steps ahead are displayed in Table 5. At these short forecast horizons BVARs and factor models outperform the AR benchmark in many cases (12, 13, and 14 cases respectively for BVAR(v.1), BVAR(v.2), and Factor model). In these instances, the forecasting gains are both large and statistically significant. In the remaining few cases in which the econometric models do not outperform the benchmark, the forecasting losses are small and never statistically significant. The TSLMs outperform the AR in fewer instances (5 for Moirai and 6 for TimesFM), also with gains that are both large and statistically significant. In the remaining cases, the forecasting losses are not statistically significant. As we have seen before, there is a tendency for TSLMs to produce the occasional more extreme results, with larger losses than those observed for the econometrics models.
$h=1$
BVAR(v.1) BVAR(v.2) Factor model BART Moirai Large TimesFM
PAYEMS 0.84 *** 0.83 *** 0.87 *** 0.97 0.83 *** 0.82 ***
INDPRO 0.90 *** 0.93 ** 0.97 1.02 0.95 ** 0.94 **
FEDFUNDS 1.21 0.97 0.87 ** 1.34 1.00 1.23
UNRATE 0.88 *** 0.84 *** 0.89 *** 0.94 1.04 1.11
RPI 0.98 0.98 0.99 1.05 1.00 1.04
DPCERA3M086SBEA 1.00 0.99 0.98 * 1.03 1.02 1.06
CMRMTSPLx 0.96 * 0.95 ** 0.98 1.01 1.01 1.04
CUMFNS 0.90 * 0.91 ** 0.95 0.96 1.09 1.48
CES0600000007 0.90 *** 0.91 *** 0.91 ** 1.05 0.93 * 0.94 **
HOUST 0.95 *** 0.95 ** 0.93 *** 1.01 1.00 0.92 ***
S&P 500 1.04 1.03 1.00 1.03 1.03 1.02
T1YFFM 1.19 1.18 1.06 1.28 1.04 1.13
T10YFFM 1.06 1.00 1.00 1.09 1.04 1.16
BAAFFM 1.02 0.96 0.94 1.07 1.01 1.28
EXUSUKx 1.02 1.03 1.01 1.05 1.02 1.01
WPSFD49207 0.97 1.00 1.03 1.03 1.01 1.01
PPICMM 1.00 1.01 0.99 1.02 1.03 1.04
PCEPI 0.94 *** 0.94 ** 0.98 1.02 0.99 0.95 *
CES0600000008 0.85 *** 0.85 *** 0.78 *** 0.91 *** 0.77 *** 0.78 ***
$h=3$
BVAR(v.1) BVAR(v.2) Factor model BART Moirai Large TimesFM
PAYEMS 0.76 ** 0.78 *** 0.79 ** 0.99 0.77 *** 0.75 **
INDPRO 0.94 0.98 0.99 0.98 0.97 0.94
FEDFUNDS 1.24 1.04 0.96 1.09 0.91 1.16
UNRATE 0.74 ** 0.72 *** 0.73 *** 0.98 0.91 0.78 **
RPI 1.00 1.00 0.99 1.02 1.00 1.01
DPCERA3M086SBEA 1.00 1.00 1.00 1.01 0.99 1.00
CMRMTSPLx 0.97 ** 0.98 0.98 1.01 1.00 1.01
CUMFNS 0.79 ** 0.85 ** 0.92 1.07 0.93 1.32
CES0600000007 0.72 *** 0.79 *** 0.79 *** 0.80 *** 0.79 *** 0.83 ***
HOUST 0.99 1.02 0.93 ** 1.04 1.01 0.78 ***
S&P 500 1.02 1.02 1.01 1.03 1.00 1.03
T1YFFM 1.32 1.30 1.14 1.32 1.03 1.12
T10YFFM 1.13 1.08 1.04 1.05 0.96 1.13
BAAFFM 1.09 1.05 1.01 1.05 0.93 1.17
EXUSUKx 1.03 1.05 1.01 1.06 1.02 1.02
WPSFD49207 0.99 1.02 1.05 1.06 0.99 1.03
PPICMM 1.00 1.02 1.00 1.05 1.00 1.03
PCEPI 0.93 ** 0.95 0.97 0.98 0.89 *** 0.91 ***
CES0600000008 0.83 *** 0.83 *** 0.77 *** 0.91 *** 0.74 *** 0.75 ***
Table 5: Forecast horizons 1 and 3 step-ahead. RMSFE relative to the AR(1). Differences in accuracy that are statistically significant at 10%, 5%, and 1% levels are denoted by one, two, or three stars, respectively. The evaluation sample is January 1985 to December 2019.
$h=6$
BVAR(v.1) BVAR(v.2) Factor model BART Moirai Large TimesFM
PAYEMS 0.99 * 0.99 * 0.99 0.81 ** 0.99 ** 0.99 *
INDPRO 0.99 1.00 1.01 1.01 0.99 0.98
FEDFUNDS 1.08 1.04 0.95 1.36 0.93 1.21
UNRATE 0.92 ** 0.92 *** 0.90 ** 0.85 0.95 0.94 **
RPI 1.00 1.00 1.00 1.03 1.00 1.01
DPCERA3M086SBEA 1.00 1.00 1.00 1.01 1.00 1.00
CMRMTSPLx 1.00 1.00 1.00 0.98 1.01 1.02
CUMFNS 0.87 * 0.90 * 1.01 1.01 0.92 1.14
CES0600000007 0.69 *** 0.76 *** 0.78 *** 0.87 *** 0.76 *** 0.79 ***
HOUST 1.01 1.05 0.94 * 0.99 0.98 0.66 **
S&P 500 1.01 1.01 1.00 1.05 1.00 1.03
T1YFFM 1.29 1.35 1.02 1.33 1.07 1.20
T10YFFM 0.98 1.05 0.92 *** 1.09 0.97 1.14
BAAFFM 1.03 1.06 0.98 1.07 0.93 1.15
EXUSUKx 1.02 1.04 1.00 1.09 1.00 1.02
WPSFD49207 1.01 1.04 1.03 1.07 0.99 1.02
PPICMM 1.02 1.04 1.00 1.04 1.01 1.06
PCEPI 0.92 0.96 0.92 1.00 0.85 *** 0.85 ***
CES0600000008 0.85 *** 0.84 *** 0.78 *** 0.83 *** 0.76 *** 0.76 ***
$h=12$
BVAR(v.1) BVAR(v.2) Factor model BART Moirai Large TimesFM
PAYEMS 1.00 0.99 1.00 1.00 1.00 0.99
INDPRO 1.01 1.01 1.01 1.04 1.01 0.99
FEDFUNDS 0.96 1.01 0.96 0.86 1.00 1.08
UNRATE 0.89 *** 0.88 *** 0.86 ** 0.99 0.96 0.89 *
RPI 1.00 1.00 1.00 1.01 1.00 1.00
DPCERA3M086SBEA 1.00 1.00 1.00 1.01 1.00 1.00
CMRMTSPLx 1.01 1.01 1.01 1.03 1.02 1.02
CUMFNS 0.91 0.93 1.06 1.08 0.92 1.02
CES0600000007 0.70 *** 0.75 *** 0.81 *** 0.80 *** 0.76 *** 0.78 ***
HOUST 1.03 1.11 1.02 1.18 0.96 0.48 **
S&P 500 1.01 1.01 1.00 1.04 1.01 1.05
T1YFFM 1.25 1.50 1.01 1.33 1.11 1.25
T10YFFM 0.87 1.05 0.90 ** 0.88 ** 1.03 1.15
BAAFFM 0.94 1.01 0.96 0.82 0.96 1.22
EXUSUKx 1.00 1.01 1.00 1.09 1.03 1.00
WPSFD49207 1.02 1.05 1.03 1.05 1.00 1.03
PPICMM 1.00 1.00 1.01 1.04 1.01 1.03
PCEPI 0.90 * 0.97 0.90 ** 0.94 0.82 *** 0.83 ***
CES0600000008 0.88 *** 0.88 *** 0.81 *** 0.91 ** 0.78 *** 0.80 ***
Table 6: Forecast horizons 6 and 12 step-ahead. RMSFE relative to the AR(1). Differences in accuracy that are statistically significant at 10%, 5%, and 1% levels are denoted by one, two, or three stars, respectively. The evaluation sample is January 1985 to December 2019.
Results for forecast horizons of 6 and 12 step-ahead are displayed in Table 6. At these horizons, for the 19 series in this medium dataset, improvements versus the benchmark are less clear-cut. In general, neither the econometric models nor the TSLMs systematically outperform the AR benchmark, even though it is worth noting that there are no instances in which they significantly under perform. Overall, Moirai is the model that more often produced better forecasts (in 13 cases at the 6-step ahead and 10 cases at the 12-step ahead) followed closely by the BVAR(v1) and factor model. For some series, there are significant gains. In particular the TSLMs produce good forecasts for Personal Consumption Expenditures (PCEPI) and Hourly Earnings (CES0600000008), with gains that are significant and larger than those obtained by the econometric models. The econometric models produce good forecasts of Unemployment rate (UNRATE) and Weekly hours (CES0600000007).
To summarize the empirical results discussed so far: i) both the econometric and the TSLM models are competitive against the AR benchmark, and never significantly outperformed by it, ii) at short horizons, the econometric models outperform the AR benchmark, while at longer horizons there is a 50:50 split between gains and losses iii) TSLMs do very well for a handful of series across all horizons.

### v4

@@ 4.1 Data
We collect 120 monthly variables for the US spanning the period January 1959 to December 2022. The data, which are obtained from the Federal Reserve Economic Data (FRED) and are available at https://fred.stlouisfed.org, cover a wide range of key macroeconomic variables that applied economists monitor regularly, such as different measures of output, prices, interest and exchange rates, and stock market performance. We provide a full list of the data and their transformations in Appendix C.
@@ 4.2 Forecasting Exercise
We use the first twenty-five year of data, January 1960–December 1984, to obtain initial parameter estimates for all the econometric models, which are then used to predict outcomes from January 1985 ($h=1$) to December 1985 ($h=12$). The next period, we include data for January 1985 in the estimation sample, and use the resulting estimates to predict the outcomes from February 1985 to January 1986. We proceed recursively in this fashion until December 2022, thus generating a time series of forecasts for each forecast horizon $h$, with h=1,,12h=1,...,12.
For the BVAR, point forecasts are obtained by taking the mean of the predictive densities. When $h>1$, the predictive densities are produced by iterating the model forward via simulation. [^16] We set the number of lags in the BVARs to $p=6$. For the factor models we first extract the first common factor, $\hat{f}_{1t}$. Next, for each variable and each forecast horizon, we run the factor-augment autoregression in (8) and forecast up to $12$ months ahead. Similarly, for BART we re-estimate the trees for each variable set and sample, as the number of series in $X_{-i,t-h}$ changes with each variable set.
We proceed in a similar fashion with all TSLMs. More specifically, for each TSLM considered, starting with the first forecast date (December 1984), we feed the historical data to the pretrained model (when working with the univariate TSLMs we supplied one time series at a time, while with the multivariate TSLMs we supplied all time series within the specific model size at once). Next, we query the model to generate a zero-shot forecast for each of the variables provided, up to 12 steps ahead. [^17] For all our experiments, we downloaded the open-sourced pretrained models and run them locally on our computer, using an Nvidia A6000 GPU with 48GB of memory. The only exception was Time-GPT, where we instead queried directly their API, with a separate query for each forecast date/time series pair. We also include a benchmark approach which uses OLS forecasts from univariate AR(1) models.
@@
where $\underline{t}$ and $\overline{t}$ denote the start and end of the out-of-sample period, and where $e_{i,j,\tau+h}^{2}$ and $e_{bcmk,j,\tau+h}^{2}$ are the squared forecast errors of variable $j$ at time $\tau$ and forecast horizon $h$ associated with model $i$ ($i\in\left\{\text{BVAR},\text{Factor model},\text{NNAR},\text{BART}\right\}\cup{TSLM}{}s$) and the AR(1) model, respectively. The point forecasts used to compute the forecast errors are obtained by averaging over the draws from the various models’ $h$-step-ahead predictive densities. Values of $RMSFE_{ijh}$ below one suggest that model $i$ produces more accurate point forecasts than the AR(1) benchmark for variable $j$ and forecast horizon $h$.
In closing this Section it is worth mentioning that in this paper we are only focusing on point forecast accuracy. This choice is intentional. While there is a rich literature on density forecasts for macroeconomic variables, using TSLMs to construct density forecasts presents additional challenges, and we leave that to further research.
@@ 5 Results
5.1 Do some TSLMs Perform Better than Others?
To begin, we focus on the overall forecasting performance of the various TSLMs under examination. Given that the evaluation sample we are considering include the Covid-19 period, to avoid conflating the overall picture with the large idiosyncrasies brought about by the pandemic, we focus first on the accuracy of the various models by stopping before the onset of the pandemic. That is, our evaluation sample stops in December 2019. Later on, we analyze how stable these results are over time, and also look at the accuracy of the various models over the most recent period, January 2020 to December 2022. Figure 2 provides a synoptic view of the forecasting performance across all 120 series we focused on, as measured by RMSFE ratios relative to the AR benchmark. Each box-plot shows the interquartile range of the distribution of relative RMSFEs for each of the TSLMs considered. The solid line within each box represents the median, while the whiskers represent the maximum and minimum relative RMSFEs that are not defined as outliers. [^18] Outliers are defined as values that are more than 1.5 times the interquantile range away from the top or bottom of the box (i.e. from the lower and upper quartiles). When the mass is concentrated above the line corresponding to $RMSFE=1$, this is an indication that the models do generally worse than the benchmark.
@@ 5.1 Do some TSLMs Perform Better than Others?
Outliers are not depicted for scaling reasons, and instead we separately report the maximum and minimum RMSFEs attained by each model in Table 2 (together with their median and standard deviation). Indeed, it does happen on several occasions that for some series the TSLMs produce forecasts that are patently unreasonable. These forecasts would be automatically disregarded by the user and would call for a fine-tuning of the models. This result highlights that while zero-shot TSLMs may produce on average reasonable forecasts, they should not be used mechanically and do require careful monitoring.
Having said that, the pattern emerging from a careful inspection of Figure 2 is one of heterogeneity among the various TSLMs. Specifically, some TSLMs show a performance that is consistently worse than others. These are TTM-enhanced and Time-GPT, for which the box plots are invariably positioned higher than those of the remaining TSLMs, signaling higher RMSFEs. These models also perform poorly compared to the benchmark, for example Time-GPT has large part of the mass of the box-plot above 1 for all horizons, and TTM for the one-step-ahead horizon. LagLlama performs a bit better than these two, however it is also systematically outperformed by the remaining models.
The TSLMs showing the best forecasting performance are Moirai and TimesFM. As we discussed, Moirai comes in three sizes, and the evidence shows that the largest size performs the best, with a mass of RMSFEs being consistently below 1 at all horizons. Because of these considerations, in the remainder of this Section, we will only focus on these best performing TSLMs (Moirai-large and TimesFM), and compare their performance against the econometric models.
5.2 Differences in Point Forecast between TSLMs and Econometric Models
@@
Focusing first on the econometric models, both the BVAR and the factor model are consistently showing a good performance relative to the AR model. This is a well known result, which has been obtained by several different studies using this dataset.
@@
As already shown in Figure 2, Moirai-large and TimesFM also perform consistently better than the AR benchmark, and Figure 3 now shows that the RMSFE distribution is similar to those obtained with the econometric models. Interestingly, while all models at almost all horizons have a median that is below the threshold of one, there is a tendency for the distribution of the RMSFEs from TSLMs to have a more pronounced right tail, signaling an increased likelihood of observing large forecast errors. Instead, econometric methods produce RMSFE distributions that are more left-skewed, signaling a consistent better performance compared to the AR(1).
Turning to the two non-linear econometric models, the evidence is mixed. The performance of BART is consistently worse than that of the benchmark. The simple NNAR model produces better forecast than BART, and is overall slightly better than the AR(1) benchmark. At 1-step ahead, the performance of NNAR is similar to that of the TSLMs. However, at longer horizons this model is outperformed by both TSLMs and the linear econometric models. This suggests that while non-linearities do play an important role, TSLMs seem to have an edge at capturing them, over the simpler non-linear models we considered.
As before, additional descriptive statistics including the maximum and minimum relative RMSFEs (including outliers) are reported in Table 3. The linear econometric models (BVARs and factor models) do not present any extreme cases in which forecast errors are a multiple of those obtained by the AR benchmark, which it is instead the case for the TSLMs results shown in Table 2, and - to a lesser extent - for the nonlinear models in the last two rows of Table 3. Moreover, it is evident that the RMFSEs coming out of the econometric models are less dispersed. To sum up, the picture is one in which BVARs and factor models seem to offer a more stable and reliable performance, when compared to zero-shot TSLMs and other simpler nonlinear models.
Note that so far we have not considered the statistical significance of the various forecasts. Appendix B presents results on the statistical significance of the various forecasts (relative to the AR(1) benchmark), focusing for ease of exposition on a subset of key macroenomic variables. These results presented in there show that for several series the improvements of both the econometric models and the TSLMs over the AR benchmark are statistically significant.